# A Circuit-QED Lattice System with Flexible Connectivity and Gapped Flat Bands for Photon-Mediated Spin Models

### v1

@@ Appendix E Flux calibration
Figure 12: Experimental setup. Schematic for the RF lines for readout and qubit control and the DC lines used to flux-bias the qubits. RF signals from a vector network analyzer (VNA) undergo $-60$ dB nominal attenuation inside the dilution fridge and pass through a $12$ GHz lowpass filter (LPF) prior to reaching the device. The output RF signals from the device pass through a second $12$ GHz LPF, then an isolator, and are then amplified by a high electron mobility transistor (HEMT) amplifier. Outside the fridge, the signals pass through a bandpass filter (BPF, $3.4$ GHz - $5.5$ GHz for measurements of the half-wave modes and $9$ GHz - $12$ GHz for measurements of the full-wave modes) and are amplified further before entering a readout port on the VNA. The DC voltage applied to each flux-bias line passes through three LPFs outside of the fridge, undergoes $-37$ dB nominal attenuation inside the fridge and passes through two more LPFs before reaching the device.
Appendix F Device Fabrication
@@
The SQUID loops were written using an Elionix ELS-G100 electron-beam lithography system. To prepare for e-beam lithography, the chip was cleaned with toluene, acetone, methanol, and isopropanol in sequence. Next, the chip was cleaned with piranha solution to remove organic residues [70]. Two layers of e-beam resist, MMA (MMA EL13) and PMMA (MMA 950 A3), were then spun onto the chip. Having two resist layers allows for the addition of undercuts in the lithography process. The sample was covered with a 30 nm anticharging layer of aluminum and then placed in the Elionix, where Manhattan-style junctions were defined through e-beam lithography. After the write, the anti-charging layer was removed with AZ 300 MIF and the sample was developed in MIBK:IPA (1:3). At this stage, the remaining e-beam resist on the chip forms channels for evaporated aluminum to fill in order to create the desired SQUID loops.
@@
Appendix G Measurement Scheme
Our quasi-1D lattice device is measured with the continuous wave (CW) measurement scheme depicted in Fig. 12. A vector network analyzer (VNA, Rohde-Schwarz ZNB20) is used to generate CW signals from up to two separated ports. The two output signals are combined with a microwave power splitter (ZFRSC-183-S+) and undergo $-60$ dB nominal attenuation inside the dilution fridge before reaching the device. The input and output of the device are protected by two identical low pass filters (LPF, K&\&&L 6L250-00089) with a cut-off frequency of $12$ GHz. An isolator (QCI-G0401202AS) on the base plate protects the device from reflected signals. The signal is then amplified by a high electron mobility transistor (HEMT) amplifier (LNF-LNC4_16B), which is specially chosen to amplify signals in the range of $4-16$ GHz encompassing both the half-wave and the full-wave modes. Finally, a post amplifier chain is located outside of the dilution fridge to amplify the signal further before it is sent into a readout port on the VNA. A band pass filter (BPF) is set before the post-amplifiers to reduce noise outside of the desired bandwidth. The frequency range chosen changes between measurements of the half-wave (3.4 GHz - 5.5 GHz) and full-wave (9 GHz - 12 GHz) modes.

### v2

@@ Appendix G Flux calibration
Figure 12: Experimental setup. Schematic for the RF lines for readout and qubit control and the DC lines used to flux-bias the qubits. RF signals from a vector network analyzer (VNA) undergo 60-60 dB nominal attenuation inside the dilution fridge and pass through a $12$ GHz lowpass filter (LPF) prior to reaching the device. The output RF signals from the device pass through a second $12$ GHz LPF, then an isolator, and are then amplified by a high electron mobility transistor (HEMT) amplifier. Outside the fridge, the signals pass through a bandpass filter (BPF, $3.4$ GHz – $5.5$ GHz for measurements of the half-wave modes and $9$ GHz – $12$ GHz for measurements of the full-wave modes) and are amplified further before entering a readout port on the VNA. The DC voltage applied to each flux-bias line passes through three LPFs outside of the fridge, undergoes 37-37 dB nominal attenuation inside the fridge and passes through two more LPFs before reaching the device.
Appendix H Device Fabrication
@@
The SQUID loops were written using an Elionix ELS-G100 electron-beam lithography system. To prepare for e-beam lithography, the chip was cleaned with toluene, acetone, methanol, and isopropanol in sequence. Next, the chip was cleaned with piranha solution to remove organic residues [Piranha_etch_deleon]. Two layers of e-beam resist, MMA (MMA EL13) and PMMA (MMA 950 A3), were then spun onto the chip. Having two resist layers allows for the addition of undercuts in the lithography process. The sample was covered with a 30 nm anticharging layer of aluminum and then placed in the Elionix, where Manhattan-style junctions were defined through e-beam lithography. After the write, the anti-charging layer was removed with AZ 300 MIF and the sample was developed in MIBK:IPA (1:3). At this stage, the remaining e-beam resist on the chip forms channels for evaporated aluminum to fill in order to create the desired SQUID loops.
@@
Appendix I Measurement Scheme
Our quasi-1D lattice device is measured with the continuous wave (CW) measurement scheme depicted in Fig. 12. A vector network analyzer (VNA, Rohde-Schwarz ZNB20) is used to generate CW signals from one or two separated ports. The two output signals are combined with a microwave power splitter (ZFRSC-183-S+) and undergo 60-60 dB nominal attenuation inside the dilution fridge before reaching the device. The input and output of the device are protected by two identical low pass filters (LPF, K&\&L 6L250-00089) with a cut-off frequency of $12$ GHz. An isolator (QCI-G0401202AS) on the base plate protects the device from reflected signals. The signal is then amplified by a high electron mobility transistor (HEMT) amplifier (LNF-LNC4_16B), which is specially chosen to amplify signals in the range of 441616 GHz encompassing both the half-wave and the full-wave modes. Finally, a post amplifier chain is located outside of the dilution fridge to amplify the signal further before it is sent into a readout port on the VNA. A band pass filter (BPF) is set before the post-amplifiers to reduce noise outside of the desired bandwidth. The frequency range chosen changes between measurements of the half-wave (3.4 GHz – 5.5 GHz) and full-wave (9 GHz – 12 GHz) modes.